parikhst:

I am trying to solve a problem and get down to w2+5w-4w-20=0 they simplified it to w(w+5)-4(w+5) cant figure out how the first came to the second equation

HeartTricker:

Hi, what exactly are we doing here?

parikhst:

well the full question is "a rectangles length is 2 units more than twice its width. its area is 40 square units. the equation w(2w+2)=40 can be used to find w, the width of the rectangle what is the width

snowflake0531:

The distributive property says \(a(b+c) = a(b) + a(c)\) and \(a(b-c) = a(b) - a(c)\) From ^ I think you can see how you came from the first equation to the second

snowflake0531:

@parikhst wrote:
well the full question is "a rectangles length is 2 units more than twice its width. its area is 40 square units. the equation w(2w+2)=40 can be used to find w, the width of the rectangle what is the width
So you have \(w(2w+2)=40\) You can distribute the parenthesis and get \(2w^2+2w=40\) then get it to one side of the equation \(2w^2+2w-40\) It's better to simplify it before finding the roots (your answer) so you can do \(2(w^2+w-20)\)

parikhst:

okay i get it to that point

snowflake0531:

From there, you can simplify it from \(2(w^2+w-20)\) to \(2(w+5)(w-4)\)

snowflake0531:

Which makes your answer either -5 or 4 But since lengths can't be negative, it must be 4

parikhst:

okay so the 2 just disapears?

snowflake0531:

Well, it's \(2(w-4)(w+5)=0\) and as you should know \(ab=c\\a=c\\b=c\) So, yeah, we just get rid of the 2

parikhst:

gotch ya okay i think that helps!

parikhst:

thanks a lot everyone

snowflake0531:

yw~

parikhst:

i also have one more question "a rectangle's width is one-fourth of its length. its area is 9 square units. the equation (1/4)*L=9 can be used to find L (length). what is length

parikhst:

i thought it would be L=36

snowflake0531:

Well, the equation is actually \[L(\frac{1}{4}L)=9\]

parikhst:

ohhh that would help!!! so u square root it duh

parikhst:

didn't see the other L

snowflake0531:

square root it-?

parikhst:

like L= square root of 36

snowflake0531:

ohhh yeah I'm slow

parikhst:

nono its ok. maybe i need my eyes checked

parikhst:

thanks!

snowflake0531:

yw~